OpenStudy (anonymous):

I need help on quadratic equations its so difficult the problem is (x+4) (x-9)=4x can someone please help me

OpenStudy (anonymous):

You need to expand the left-hand side and collect like terms. You will end up with another quadratic which you can factor and solve.

OpenStudy (anonymous):

\[(x+4)(x-9)=4x \iff x^2+4x-9x-36=4x \iff x^2-9x-36=0\]which may be factored\[(x+3)(x-12)=0\]

OpenStudy (anonymous):

x^2-9 x-36 = 0

OpenStudy (anonymous):

Your solutions are then \[x=-3, x=12\]

OpenStudy (anonymous):

(x-12)(x+3)

OpenStudy (anonymous):

Foil the (x+4)(x-9) which comes out as x^2-5x-36=4x then you subtract 4x on both sides which leaves you with x^2-9x-36=0 then factor it to (x-12)(x+3)=0 x=12 x=-3

OpenStudy (anonymous):

Thank you all very much

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Can someone explain to me how to do quadratic equations, im confused?

OpenStudy (anonymous):

Look under 'algebra' on this site. In the second column, there should be stuff for the quadratic equation. http://www.khanacademy.org/